Making Protest Signs

Here is how we made the signs we used in the Lotus demonstration.

1. Get corrugated cardboard approximately 2 ft. by 3 ft. One way to get it is to cut up large boxes. You can also get it from an art supply store.

2. Get sticks to hold the signs with. We used sticks 0.75 in. thick and 1.5 in. wide; sticks much thinner may not hold the weight.

3. Paste white paper onto one side, and draw the sign on it. Or make large letters with a computer printer and paste them on above the background paper.

4. Staple through the cardboard into a wooden stick with a heavy duty stapler. Then tape the stick onto the cardboard in back with plastic packing tape.

5. If you can get access to a shrink wrap machine, it is good to shrink wrap the sign to protect it from rain. Otherwise, if you want to be able to risk bad weather, tape the whole front of the sign with more plastic packing tape. That will do for light rain.

6. Use plastic packing tape to wrap the stick so that it doesn’t dig into the skin or give splinters.

People have fun drawing their own signs or inventing slogans. However, here are a few to start with:

Innovate, don’t litigate.

Boycott Lotus and Apple

Users want compatible software.

Drop the suit, we have you surrounded.

Don’t make me wear your suit (with picture of straightjacket).

League for Programming Freedom

A large image of Liberty Empowering the Programmers also makes a good sign. We can send you them if you need them.